• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Saha

    Log entry time 12:11:57 on December 4,2001

    Entry number 78723

    keyword=Hall OTR update from J.C. Denard

    Hall A target OTR. Foil changed - J.C. Denard (ELOG #165989)

    Karel Capek and I changed the hall A target carbon foil which was damaged. A new foil identical to the old one was ready to install and has been installed. However, looking at the old foil, we saw it was actually damaged by the beam. This should not have happened if the thermal conductivity of the material had been as expected.
    Then we suspect the new foil will probably deteriorate also. Karel is planning to mount a better foil this week in order to install it next week after hall A goes into a down for a new experiment. We will know for sure in January if the better foil is OK.